A coperation is trying to decide whether or not to buy the patent for a product designed by another company. The decision to buy means of investment of 8 million...A corporation is trying to decide whether or not to buy the patent for a product designed by another company. The decision to
Answer #1

1) $11802088

since the investment is $8 and the outcomes are different for different periods of time interest given 10%

when the demand is light

8*10%*3 = 24x/10 = 1.3

x1 = 13/24

when the demand is moderate

8*10%*4 = 32x/10 = 2.5

x2 = 25/32

when the demand is high

8*10%*4 = 32x/10 = 4

x3 = 40/32

Expectation = summation(xi * probability)

= x1* 0.2 + x2 * 0.6 + x3 * 0.4

= 11802088 dollars

2) standard deviation = $217669

3) coefficient of variation = 217669 / 11802088

= $18443